OpenStudy (anonymous):

1. A spinner numbered 1 through 5 is spun 3 times. What is the probability of spinning a number less than 3, then a number greater than 3, and then the number 3? Write your answer as a decimal. Do not round. 2. A bag contains 8 red, 4 blue, and 4 yellow marbles. What is the probability of randomly choosing 2 blue marbles without replacement? Write your answer as a decimal. Round to the nearest hundredth if needed.

OpenStudy (triciaal):

!) less than 3 2/5 greater than 3 2/5 3 = 1/5 probability of and you multiply 2*5*2/5*1/5 = 4/125 as a decimal = 0.032

OpenStudy (xapproachesinfinity):

the question didn't and the probability of getting greater than 3,less than 2, and 3 it simply asked to find each individual prob

OpenStudy (xapproachesinfinity):

ask*

OpenStudy (xapproachesinfinity):

so no need to multiply anything

OpenStudy (triciaal):

is spun 3 times What is the probability of spinning a number then and then I read it twice myself before deciding it was combined

OpenStudy (xapproachesinfinity):

eh, right! i kinda happen to forgot about how many spinning